### v1

# An Introduction to Zero-Order Optimization Techniques for Robotics

Abstract
Zero-order optimization techniques are becoming increasingly popular in robotics due to their ability to handle non-differentiable functions and escape local minima. These advantages make them particularly useful for trajectory optimization and policy optimization. In this work, we propose a mathematical tutorial on random search. It offers a simple and unifying perspective for understanding a wide range of algorithms commonly used in robotics. Leveraging this viewpoint, we classify many trajectory optimization methods under a common framework and derive novel competitive RL algorithms.
I Introduction
In recent years, zero-order (or derivative-free) optimization techniques [1] have gained a lot of popularity in the robotics community. While zero-order optimization is a well-established field, its widespread deployment in robotics has only been made possible by the recent advances in parallel computing and GPU hardware. These improvements have enabled the use of zero-order Trajectory Optimization (TO) techniques in a Model Predictive Control (MPC) fashion on complex robots [2, 3]. In parallel, Reinforcement Learning (RL) has emerged as a powerful tool and demonstrated state-of-the-art capabilities in tasks such as locomotion or manipulation [4, 5].
While these approaches may appear unrelated at first, they share a common property: they do not require access to the simulator’s gradients. This characteristic allows them to optimize non-smooth objective functions, which typically arise in problems involving contact such as locomotion or manipulation. Another benefit is that it avoids the tedious development of efficient gradient implementations. Furthermore, while deterministic gradient-based algorithms, such as gradient descent or Newton method, are prone to getting stuck in local minima, zero-order techniques are mostly stochastic, which can help escape local minima.
Zero-order optimization has a long history [6]. In its most generic form, zero-order optimization performs random search by generating samples at random and updating an estimate of the optimal solution based on the performance of each sample. In the 1970s, Evolutionary Strategies (ES) gained popularity [7] and later culminated with the CMA-ES algorithm [8]. In TO, predictive sampling [9] is a well-known and simple random search technique. Another popular algorithm is MPPI [10], which was initially derived from information theory. In RL, Reinforce [11] was one of the first major algorithms proposed for continuous action space problems. These algorithms have since inspired many variations.
While most of these algorithms have strong theoretical foundations, this is not necessarily well known in the robotics community. One reason for this is that these algorithms often fall at the intersection of many fields such as optimization, statistics, machine learning and control. By looking at results from these different fields, we aim to answer two questions: 1) What are the convergence guarantees of the algorithms used in robotics? 2) What are the theoretical tools explaining the ability of these algorithms to escape local minima?
Recently, [12] introduced Gaussian smoothing and understood random search from an optimization perspective. Inspired by this work, we propose a unified perspective to understand the zero-order optimization techniques popular in the robotics community. More specifically, we are interested in algorithms that can find solutions to the following problem:
$\min_{x\in\mathbb{R}^{n}}f(x)$ (1)
This formulation encompasses a wide range of problems encountered in robotics, such as TO and RL. Both of these techniques define a high-level goal through a cost function that has to be minimized (or a reward function that has to be maximized). However, they differ in the optimization space. TO optimizes directly over trajectories, whereas RL optimizes over policies. Furthermore, they differ in the way they are deployed. These algorithms can be used either online (i.e. executed at runtime) or offline. On one hand, TO can either be solved online in an MPC loop or solved offline to generate data to train a policy. On the other hand, RL is mostly solved offline. While these differences may seem fundamental, we show that the algorithms used to solve these problems share many conceptual similarities.
Related work
[13] first showed the similarities between MPPI and CMA-ES. More recently, [14] understood MPPI as an approximate gradient method and [3] connected MPPI to diffusion models. Our work builds on these insights and further investigates explanations for the success of MPPI. In particular, we shed light on the connection between MPPI and recent results showcasing the benefits of the log-sum-exp transform [15]. Lastly, we benchmark several state-of-the-art approaches on various robotic examples.
[16, 17] first made the connection between RL and random search. These works used this insight to derive TO algorithms based on random search. In contrast, we study how to leverage this understanding to derive novel RL algorithms.
Contribution
In this work, we propose a mathematical introduction to zero-order optimization algorithms used in robotics. This unified perspective provides a simple way to understand techniques such as MPPI, Covariance Matrix Adaptation (CMA) or RL policy gradient methods. In addition, we show how this unified view allows us to naturally derive novel competitive methods. Lastly, we discuss the theoretical concepts explaining why stochastic algorithms are well-suited to avoid getting stuck in local minima.
Outline
Section II introduces the main concepts of random search. Section III connects popular TO algorithms to random search. Section IV investigates the connections between Gaussian smoothing and RL. Lastly, Section V discusses how to improve the sample efficiency of these algorithms by relying on a population of samples.
II Random search
II-A Simple random search
In this first section, we present how to use a population of samples (or particles) to search for a global solution. We divide the population search algorithms in two broad classes: global search and local search algorithms.
II-A1 Global Search
Global search can be considered as the simplest global optimization technique. The idea is to iteratively sample a random point. If the sample has a value lower than the previous one, it is kept as the current best guess. Otherwise, it is disregarded. Algorithm 1 summarizes the procedure.
Input: $x_{0}\in\mathbb{R}^{n}$
1 $x\leftarrow x_{0}$
2 while stopping criterion is not met do
3       Sample $\tilde{x}$ in $\mathbb{R}^{n}$
4       if $f(\tilde{x})<f(x)$  then
5             $x\leftarrow\tilde{x}$
6
7
Output: $x$
Algorithm 1 Simple Global Search
Interestingly, the sampling distribution of Algorithm 1 ignores all the previous estimates. For this reason, it can be referred to as a global algorithm. Under some mild conditions on the objective function, Algorithm 1 converges to a global minimum [18]. Unfortunately, in practice, this method performs poorly as it is subject to the curse of dimensionality. That is to say that the number of function evaluations required to find a global solution grows exponentially with the dimension of the problem.
II-A2 Greedy Local Search
A common attempt to break the curse of dimensionality is to exploit problem-specific knowledge. For instance, exploiting previous information about the shape of the function can help design an empirically more efficient algorithm. This is called Greedy Local Search. The idea is to iteratively search around the previous estimate, with additive Gaussian noise, for instance. Then, as in the previous algorithm, the sample is kept only if it improves the function value. Algorithm 2 summarizes the procedure.
Input: $x_{0}\in\mathbb{R}^{n}$
1 $x\leftarrow x_{0}$
2 while stopping criterion is not met do
       Sample $d$ in $\mathbb{R}^{n}$ // with a centered Gaussian distribution
3       if $f(x+d)<f(x)$  then
4             $x\leftarrow x+d$
5
6
Output: $x$
Algorithm 2 Greedy Local Search
By design, Algorithm 2 depends on previous estimates. For that reason, it can be referred to as a local algorithm. However, this notion should not be confused with the concepts of local and global solutions. In fact, although the convergence guarantees of Algorithm 2 are more restrictive than those of Algorithm 1, it is still possible to prove convergence of Algorithm 2 to a global solution under some regularity conditions [18]. Consequently, both local and global search can have global convergence guarantees. However, while both are subject to the curse of dimensionality, local search techniques can be more effective empirically. In fact, as we will see in the next section, it is well-known that local search techniques relying on gradient information can exhibit strong convergence rates and guarantees in convex settings.
II-B Random search via gradient approximation
In this section, we present a few ways to perform approximate gradient descent using only function evaluations. The class of methods under study is summarized in Algorithm 3.
Input: $x_{0}\in\mathbb{R}^{n}$
1 $x\leftarrow x_{0}$
2 while stopping criterion is not met do
       Sample $g$ approximating the gradient at $x$. // e.g. Eq. (3), Eq. (4), Eq (9) or Eq (10)
3       $x\leftarrow x-\alpha g$
Output: $x$
Algorithm 3 Approximate Gradient Descent
Here, $\alpha$ is the step size (or learning rate). There exist many ways to estimate the gradient with multiple function evaluations [19]. Among the most widely used ones is the forward finite differences technique:
$\sum_{j=1}^{n}\frac{f(x+\mu e_{j})-f(x)}{\mu}e_{j}$ (2)
where $e_{j}\in\mathbb{R}^{n}$ represents the $j^{th}$ canonical vector. Alternatively, one may use central finite differences. While this estimate can be very effective, it requires $n+1$ function evaluations, which is a limitation in high-dimensional settings.
II-B1 Random Coordinate Descent
One way to avoid performing that many function evaluations is Random Coordinate Descent [20]. As its name suggests, this method alternates descent steps on random coordinates.
$g=\frac{f(x+\mu e_{j})-f(x)}{\mu}e_{j}$ (3)
where $j$ can be chosen uniformly at random in $\{1,\dots,n\}$. Note that $g$ depends on $x$.
Remark 1.
In the convex case and without finite difference approximation, Random Coordinate Descent requires at most $n$ times more steps compared to classical gradient descent [20]. In addition, on a certain class of functions, it can require as many steps as gradient descent while requiring $n$ times less function evaluations.
Random Coordinate Descent samples discrete coordinates, but it is also possible to sample along arbitrary directions.
II-B2 Simultaneous Perturbation Stochastic Approximation (SPSA)
In the 1990s, Spall introduced Simultaneous Perturbation Stochastic Approximation (SPSA) [18]. The main idea is to replace finite differences with the following stochastic estimate:
$g=\frac{f(x+\mu\Delta)-f(x-\mu\Delta)}{2\mu}\Delta$ (4)
where each component of $\Delta$ is equal to ±$\pm 1$ according to independent Bernoulli distributions. Note that Spall derived SPSA using central differentiation instead of the forward finite difference. Under reasonable assumptions, SPSA can reach similar performances as finite differences while requiring $n$ times less samples [18].
II-B3 Gaussian Smoothing
In [12], Nesterov proposes another approach that samples the gradient of a smoothed surrogate function
$f_{\mu}(x)=\mathbb{E}[f(x+\mu u)]$ (5)
where $u\sim\mathcal{N}(0,\Sigma)$. This is Gaussian smoothing, commonly referred to as randomized smoothing. As $\mu$ goes to zero, $f_{\mu}$ tends to $f$. As a result, for small values of $\mu$, the gradients of the surrogate function $f_{\mu}$ are close to those of the original function $f$. The advantage of this surrogate is that its gradient can be evaluated via an average of function evaluations. To understand this, let’s first define $\kappa=\sqrt{(2\pi)^{n}\det(\Sigma)}$, the normalization factor of the multivariate Gaussian distribution. With a change of variables, we can write:
[…]
Consequently,
[…]
$=\mathbb{E}\left[\frac{1}{\mu}f(x+\mu u)\Sigma^{-1}u\right]$
$=\mathbb{E}\left[\frac{f(x+\mu u)-f(x)}{\mu}\Sigma^{-1}u\right]$ (7)
The last equality comes from the fact that $\mathbb{E}\left[f(x)\Sigma^{-1}u\right]=0$. Hence, the gradient of $f_{\mu}$ can be estimated using only function evaluations of the original function, $f$ (at the cost of approximating an expectation with samples). Note that the term $\Sigma^{-1}u$ is the gradient of the logarithm of the probability density function with respect to the mean; this is often called the log-likelihood trick. This is sometimes written:
$\nabla f_{\mu}(x)=\mathbb{E}_{z\sim\mathcal{N}(x,\Sigma)}\left[f(z)\nabla_{x}\log(p_{x}(z))\right]$ (8)
where $p_{x}$ is the density of the Gaussian distribution with mean, $x$, and with covariance, $\Sigma$. This alternative view will be useful to understand CMA. In the general case, the expectation is intractable, however, one can rely on sampling to estimate it. For instance, one can use the following estimate in Algorithm 3:
$g=\frac{f(x+\mu u)-f(x)}{\mu}\Sigma^{-1}u$ (9)
While $\frac{1}{\mu}f(x+\mu u)\Sigma^{-1}u$ is a valid gradient estimate, it can be arbitrarily large. In contrast, the estimate (9) is invariant to constant translations of the function. Intuitively, this will help reduce the variance of the estimate. Note that it is also possible to use an estimate with central differentiation:
$g=\frac{f(x+\mu u)-f(x-\mu u)}{2\mu}\Sigma^{-1}u$ (10)
In the convex setting, using the forward estimate (Eq. (9)) or the central estimate (Eq. (10)) as a gradient estimate requires at most $n$ times more iterations than the standard gradient method [12].
Remark 2.
Alternatively, one may sample from the surface of a unit sphere [21]. This technique guarantees that the norm of the direction vectors is bounded.
To summarize, although the assumptions in [18, 20, 12] vary, the underlying ideas are of the same nature; sampling along random directions may match the performance of gradient descent while requiring $n$ times less function evaluation.
II-C Simulated Annealing
So far, we have seen that performing random approximate gradient descent can be sample efficient in the convex setting. However, one may ask whether those algorithms can converge to global solutions in nonconvex settings, as it is the case for the Greedy Local Search introduced in Algorithm 2. It is well known that, in its vanilla form, gradient descent is prone to local minima. However, additive Gaussian noise can lead to asymptotic global convergence guarantees. To understand why, we need to study the following Langevin dynamics:
$dX_{t}$ $=\nabla_{x}\log p(X_{t})+dW_{t}$ (11)
Here $p$ is a probability distribution and $W_{t}$ is a Brownian motion. $\nabla_{x}\log p$ is the score function of the distribution. Note that the gradient is taken with respect to the input of $p$ (not with respect to parameters of $p$ as in the log-likelihood trick of Gaussian smoothing). This Stochastic Differential Equation (SDE) has been extensively studied and it can be shown that $X_{t}$ converges to the distribution defined by $p$ [22]. Consequently, this SDE can be used to sample from a distribution with probability density $p$. One way to leverage this result in the context of global optimization is to choose the following probability distribution:
$p\propto\exp\left(-\frac{1}{\lambda}f\right)$ (12)
where $\lambda>0$ is called the temperature. As $\lambda\rightarrow 0$, the distribution concentrates on the set of global minima of $f$ [23]. Furthermore, the score function of this distribution is proportional to the gradient of $f$:
$\nabla_{x}\log p(x)=-\frac{1}{\lambda}\nabla_{x}f(x)$ (13)
Therefore, this SDE can be discretized in order to derive a recursive algorithm. Following this idea, [24] shows that with an appropriate schedule of step sizes $\alpha_{k}$ and $\gamma_{k}$, the following discretized Stochastic Langevin Gradient Dynamics guarantees asymptotic convergence to global solutions.
$x_{k+1}=x_{k}-\alpha_{k}g_{k}+\gamma_{k}w_{k}$ (14)
where $w_{k}$ is Gaussian noise and where $g_{k}$ is of the form $(\nabla f(x_{k})+\xi_{k})$ with $\xi_{k}$ representing stochastic noise. Hence, $g_{k}$ can be interpreted as a noisy approximation of the gradient. Intuitively, the Gaussian noise $w_{k}$ helps to escape local minima. The step sizes $\alpha_{k}$ and $\gamma_{k}$ are chosen such that the corresponding SDE has a temperature $\lambda$ that tends slowly to zero. This is called Simulated Annealing. Intuitively, slowly decreasing the temperature will guide the iterates towards a global solution.
These results can be used to prove the global convergence of SPSA. More precisely, the iterative scheme in Equation (14) with $g$ as the SPSA estimate defined in Equation (4) converges to a global minimum [25]. In fact, in its default version (i.e. without additive noise), SPSA converges to a global minimum [25]. This is because the noise of the estimate (i.e. $\xi_{k}$) is enough to escape local minima and to achieve global convergence. We are not aware of such results in the context of randomized smoothing. However, we can conjecture that similar results may be valid.
To summarize, we have seen that stochasticity can help gradient descent algorithms to reach global solutions. Although these results are highly relevant theoretically, they have limited practical impact as they are only asymptotic.
Remark 3.
At first sight, it might be hard to relate the discretized Langevin update to the Greedy Local Search in Algorithm 2. However, following the observation from [26], the if statement can be related to the Metropolis-Hastings algorithm [23]. We refer the reader to the Appendix A for more details.
III Trajectory Optimization
In this section, we show how random search allows us to understand derivative-free optimization algorithms commonly used in robotics. In practice, Trajectory Optimization aims to solve
$\min_{\{u_{0},u_{1},\dots u_{T-1}\}}$ $\;\sum_{t=0}^{T-1}c_{t}(x_{t},u_{t})+c_{T}(x_{T})$ (15)
such that $\,x_{0}=x,\,\mbox{and}\;\;x_{t+1}=f_{\text{dyn}}(x_{t},u_{t})$
Here $x$ denotes the state, $u$ the control variable, $(c_{t})_{t}$ the running cost functions, $c_{T}$ the terminal cost function, $T$ the time horizon and $f_{\text{dyn}}$ the dynamics. Note that we abuse the notation by using $x$ to denote the state instead of the optimization variable. Importantly, the constraints are implicit; the optimization is only done with respect to the control variables. Therefore, this problem is unconstrained and called the single shooting approach. In the end, it can be written in the form of Problem (1).
III-A Predictive sampling
In robotics, Predictive sampling [9] is one of the most simple random search approach. This method can be seen as a variation of the Greedy Local Search (Algorithm 2). Instead of sampling one search direction, it samples $K$ search directions and keeps the best. If no direction provides an improvement, the previous estimate is kept. Algorithm 4 summarizes the procedure.
Input: $x_{0}\in\mathbb{R}^{n}$, K
1 $x\leftarrow x_{0}$
2 while stopping criterion is not met do
       Sample $d_{1},\dots,d_{K}$ // with independent centered Gaussian distributions
3       $D\leftarrow\{x\}$
4       for $k=1,\dots K$ do
5             Add $x+d_{k}$ to $D$
6      $x\leftarrow\operatorname{argmin}_{\tilde{x}\in D}f(\tilde{x})$
Output: $x$
Algorithm 4 Predictive Sampling
Importantly, all the function evaluations can be performed in parallel. While this approach can be quite effective, recent MPC demonstrations rely on more sophisticated algorithms that use gradient information.
III-B The log-sum-exp transform: MPPI
Model Predictive Path Integral (MPPI) [10] is a derivative-free TO method initially derived from an information theoretic point of view. MPPI iteratively samples around the current guess and updates it using a weighted exponential average. More specifically, given a current guess $x$, MPPI samples $K$ variables $x_{k}$ following a Gaussian distribution, $\mathcal{N}(x,\Sigma)$. Then, it takes an update according to the following rule:
$x\leftarrow\sum_{k=1}^{K}w_{k}x_{k}$ (16)
where
$w_{k}=\frac{\exp(-\frac{1}{\lambda}(f(x_{k})-\rho))}{\sum_{j=1}^{K}\exp(-\frac{1}{\lambda}(f(x_{j})-\rho))}$ (17)
where $\rho=\min_{j}f(x_{j})$ (to prevent numerical stability issues). These weights, $w_{k}$, are called Exponential Average weights. Algorithm 5 describes the complete procedure.
Input: $x_{0}\in\mathbb{R}^{n},\Sigma,K$
1 $x\leftarrow x_{0}$
2 while stopping criterion is not met do
       Sample $x_{1:K}\sim\mathcal{N}(x,\Sigma)$
        // sample
3       Compute weights $w_{1},\dots w_{K}$
       $x\leftarrow\sum\limits_{k=1}^{K}w_{k}x_{k}$
        // update
4
Output: $x$
Algorithm 5 Model-Predictive Path Integral (MPPI)
Let’s see how this update rule can be interpreted as an instance of Gaussian smoothing. Instead of the surrogate function defined in Equation 5, let’s consider the (continuous) log-sum-exp function.
$f_{\mu,\lambda}(x)=-\lambda\log\left(\mathbb{E}\left[\exp\left(-\frac{1}{\lambda}f(x+\mu\epsilon)\right)\right]\right)~{}$ (18)
where $\epsilon\sim\mathcal{N}(0,\Sigma)$ and where $\lambda>0$ is called the temperature. As $\mu$ goes to zero, we recover $f$. Therefore, the gradients of this surrogate can be used to approximate the gradients of the original function, $f$. Furthermore, as $\lambda\rightarrow\infty,f_{\mu,\lambda}\rightarrow f_{\mu}$ [15]. Consequently, this surrogate can be seen as a generalization of Gaussian smoothing. The gradient of the surrogate reads:
$\nabla f_{\mu,\lambda}(x)=\frac{-\lambda\mathbb{E}\left[\exp\left(-\frac{1}{\lambda}f(x+\mu\epsilon)\right)\Sigma^{-1}\epsilon\right]}{\mu\mathbb{E}\left[\exp\left(-\frac{1}{\lambda}f(x+\mu\epsilon)\right)\right]}~{}$ (19)
In practice, both expectations can be approximated with $K$ samples $\epsilon_{1:K}$ following the distribution $\mathcal{N}(0,\Sigma)$. This leads to the following estimate:
$g=\frac{-\lambda}{\mu}\frac{\sum\limits_{k=1}^{K}\exp\left(-\frac{1}{\lambda}f(x+\mu\epsilon_{k})\right)\Sigma^{-1}\epsilon_{k}}{\sum\limits_{j=1}^{K}\exp\left(-\frac{1}{\lambda}f(x+\mu\epsilon_{j})\right)}$ (20)
Similarly to the randomized smoothing case, this update is invariant to additive translation of the function $f$. Let’s consider $\mu=1$. We will show that MPPI follows a natural gradient step [27]:
$x\leftarrow x-\alpha F^{-1}g$ (21)
where $\alpha$ is the step size and where $F$ is the Fisher information matrix which is equal to the inverse of the covariance matrix for Gaussian distributions. The idea of the natural gradient is to build an update step that is invariant to changes of variables. More details on the natural gradient are provided in the appendix B. If $\alpha=\frac{1}{\lambda}$, we have
$x-\alpha F^{-1}g$ $=x+\frac{\sum\limits_{k=1}^{K}\exp\left(-\frac{1}{\lambda}f(x+\mu\epsilon_{k})\right)\epsilon_{k}}{\sum\limits_{j=1}^{K}\exp\left(-\frac{1}{\lambda}f(x+\mu\epsilon_{j})\right)}$
$=\sum\limits_{i=1}^{K}w_{k}(x+\epsilon_{k})$ (22)
As $(x+\epsilon_{k})\sim\mathcal{N}(x,\Sigma)$, we recover the MPPI update. Furthermore, if $\Sigma=\sigma^{2}I$, then $\frac{\sigma^{2}}{\lambda}\leq\frac{1}{L}$ where $L$ is the Lipschitz function of the surrogate function [15]. Therefore, the step size chosen by MPPI can be interpreted as a conservative estimate of the standard optimal step size from convex optimization [28].
Figure 1 illustrates how the log-sum-exp smoothing differs from Nesterov’s randomized smoothing. This simple example function has two types of local minima: one minima is global but narrow and the other one is local but wide. Interestingly, Gaussian smoothing gives similar importance to both minima; the two minima now have a similar depth and width. In contrast, the log-sum-exp smoothing gives more importance to the global minima; the global minima is much wider and almost absorbs the local one.
Figure 1: randomized smoothing (RS) compared to log-sum-exp (LSE) smoothing with $\lambda=0.1$.
To the best of our knowledge, [15] proposes the first explanation of the success of MPPI from an optimization perspective. [15] shows that the log-sum-exp smoothing can be interpreted as an interpolation between the default randomized smoothing and the Moreau envelope. The idea is that the Moreau envelope is a better approximation than RS but is harder to compute. MPPI elegantly trades-off approximation quality and computational complexity. At the cost of more samples, relying on this surrogate can improve convergence rates [15]. Intuitively, this estimator requires many samples because of the presence of the expectation in the denominator in Equation (19). With the progress of parallel computing, this requirement of multiple samples might not be a limitation at all. However, using many samples will render the update step essentially deterministic, which will make the algorithm prone to getting trapped in local minima.
Remark 4.
The log-sum-exp smoothing reminds of the risk-seeking control formulation [29]. We refer the reader to Appendix C for a discussion on the topic.
Remark 5.
The log-sum-exp smoothing reminds of the simulated annealing idea. Similarly to simulated annealing, as the temperature $\lambda$ goes to 00, the surrogate concentrates around the global minima. In fact,
$p\propto\exp\left(-\frac{1}{\lambda}f_{\mu,\lambda}\right)=\mathbb{E}\left[\exp\left(-\frac{1}{\lambda}f(x+\mu\epsilon)\right)\right]$
Therefore, applying the Langevin update to the log-sum-exp surrogate function amounts to applying Gaussian smoothing to the probability distribution proportional to $e^{-\frac{1}{\lambda}f}$.
III-C Covariance Matrix Adaptation (CMA)
So far, we have kept the covariance matrix $\Sigma$ fixed across various iterations. However, in practice, it is not necessarily clear how to choose this parameter according to the problem. In this section, we show how to derive a Covariance Matrix Adaptation (CMA) scheme. Interestingly, the optimization of the smooth surrogate function can be interpreted as an optimization in the space of Gaussian probability distribution. Indeed, optimizing Equation (5) can be interpreted as a search over the mean of a Gaussian distribution. More specifically, we can write
$f_{\mu}(x)=\mathbb{E}[f(x+\mu u)]=\mathbb{E}_{z\sim\mathcal{N}(x,\Sigma)}\left[f(z)\right]$ (23)
Therefore, Gaussian smoothing can be seen as the search of a Gaussian distribution that minimizes the expectation of $f$ under that distribution. This distribution can be interpreted as the belief of where the global minimum of the function $f$ might be. A natural generalization is to not only optimize the mean but also the covariance. More specifically, one can apply gradient descent to the following problem:
$\min_{\theta=(x,\Sigma)}J(\theta)$ (24)
where
$J(\theta)=\mathbb{E}_{z\sim\mathcal{N}(x,\Sigma)}\left[f(z)\right]$ (25)
$J(\theta)$ is minimal when the distribution $\mathcal{N}(x,\Sigma)$ concentrates around the minima of $f$ [30]. Note that this idea do not necessarily need to rely on Gaussian distributions. Similarly to the randomized smoothing case, we can use the log-likelihood trick:
$\nabla J(\theta)=\mathbb{E}_{z\sim\mathcal{N}(x,\Sigma)}\left[f(z)\nabla_{\theta}\log(p_{\theta}(z))\right]$ (26)
Here, the gradient of the probability distribution is taken with respect to both the mean and the covariance. Akimoto et al. [31] shows that the natural gradient update $\Delta\theta=F(\theta)^{-1}\nabla J(\theta)$ can be written:
$\Delta\Sigma$ $=\mathbb{E}_{z\sim\mathcal{N}(x,\Sigma)}\left[f(z)(\left(z-x\right)\left(z-x\right)^{\top}-\Sigma)\right].$
$\Delta m$ $=\mathbb{E}_{z\sim\mathcal{N}(x,\Sigma)}\left[f(z)(z-x)\right]$ (27)
Similarly to the randomized smoothing case, this update step averages quantities that only rely on function evaluation of the original function. Following the natural gradient is especially relevant in this setting, as it is crucial that the update does not depend on the parametrization choice of the Gaussian distribution. We refer the reader to the appendix B for more details on the natural gradient. In practice, this update step can be approximated by samples $x_{k}\sim\mathcal{N}(x,\Sigma)$. Denoting $w_{k}=f(x_{k})$, the approximate update with a step size $\alpha$ reads
$\Sigma$ $\leftarrow(1-\alpha\sum\limits_{k=1}^{K}w_{k})\Sigma+\alpha\sum\limits_{k=1}^{K}w_{k}(x_{k}-x)(x_{k}-x)^{\top}$
$x$ $\leftarrow(1-\alpha\sum\limits_{k=1}^{K}w_{k})x+\alpha\sum\limits_{k=1}^{K}w_{k}x_{k}$ (28)
This allows us to recovers the CMA scheme (see Algorithm 6). Note that it is important to update the covariance matrix first as the covariance update must rely on the old $x$ (i.e. the one used to sample). As with Gaussian smoothing, a constant can be subtracted to $f$ without changing the expectation in Equation (26). Therefore, another choice of weights could be $w_{k}=f(x_{k})-f(x)$ as this renders the update invariant to translations of $f$. One could go further and try to design an algorithm invariant to any increasing transformation of the objective function. This can be done by replacing $f(x_{k})$ by ordered weights $w_{1}\leq w_{2}\leq\dots\leq w_{K}$. As pointed out by [31, 32, 30], this recovers the CMA-ES algorithm (without evolution path) [8]. [30] provides a theoretical justification for this choice of weights.
Input: $x_{0}\in\mathbb{R}^{n},\Sigma_{0},K,\alpha$
1 $(x,\Sigma)\leftarrow(x_{0},\Sigma_{0})$
2 while stopping criterion is not met do
3       Sample $x_{1:K}\sim\mathcal{N}(x,\Sigma)$
4       Compute weights $w_{1},\dots w_{K}$
5       $\Sigma\leftarrow(1-\alpha\sum\limits_{k=1}^{K}w_{k})\Sigma$ $\phantom{.......}+\,\,\alpha\sum\limits_{k=1}^{K}w_{k}(x_{k}-x)(x_{k}-x)^{\top}$
6       $x\leftarrow(1-\alpha\sum\limits_{k=1}^{K}w_{k})x+\alpha\sum\limits_{k=1}^{K}w_{k}x_{k}$
Output: $x$
Algorithm 6 Covariance Matrix Adaptation (CMA)
Remark 6.
Another choice of weight could be to assign $w_{k}=\frac{1}{K_{e}}$ to the best $K_{e}$ samples and 00 to the others (by best, we mean samples with the lowest value $f(x_{k})$). As pointed out by [30], this elitist weighting provides an interesting connection to the well-known Cross-Entropy Method (CEM). We refer the reader to the appendix D for more details.
An important point that we overlooked is whether the covariance update rule in Equation (III-C) will maintain the matrix positive definite. If $10\leq\eta<$ 1, $w_{i}\geq 0$ and the weights sum to one, then the covariance matrix will always be positive definite provided that we started with a positive definite matrix [31]. However, in practice, there is no reason that the function $f$ would be such that the weights satisfy this property.
One possible solution is to design an update that operates on exponential coordinates so that the covariance matrix naturally stays positive definite. This is the idea behind xNES [32]. Thanks to the invariant properties of the natural gradient, both update rules align when the step size goes to zero.
Another approach could be to use the log-sum-exp transformation used for MPPI. Indeed, this transformation naturally maintains the weights positive and such that they sum to one. We omit the derivations as it is straightforward to show that applying the natural gradient to Equation (18) yields the update rule (III-C) with the Exponential Average weights (Eq (17)). We will refer to this algorithm as MPPI-CMA. Table I summarizes all the possible variations of the CMA algorithm depending on the choice of weight.
Weights
MPPI-CMA Exponential Average (Eq 17)
CMA-ES Ordering [8]
CEM⋆⋆ Elitist (Rmk 6)
TABLE I: Variations of algorithm 6. We refer to CMA-ES without evolutionary path. ⋆⋆Note that CEM updates the mean first (see Appendix D).
When dealing with TO, it is often preferred to design algorithms that can achieve linear complexity with respect to the time horizon $T$ (Equation (15)). This is typically achieved by exploiting the problem’s structure. One way to do so is to use a block-diagonal covariance matrix (with $T$ blocks of size $n_{u}$ where $n_{u}$ is the dimension of the control input), i.e. $\Sigma=\operatorname{blockdiag}(\Sigma_{1},\dots\Sigma_{T})$. Clearly, this makes the update rules of Predictive Sampling and MPPI easily implementable with linear complexity with respect to the time horizon.
However, this is less straightforward in the case of the CMA update. Specifically, Equation (III-C) will not necessarily maintain the block-diagonal structure of the covariance matrix. This is because $\Sigma$ is not the appropriate representation to work with. Instead of applying the natural gradient descent to the full matrix $\Sigma$, one can apply it to each block $\Sigma_{t}$. By construction, this approach will maintain the overall covariance matrix block-diagonal. Intuitively, the natural gradient update for each block matches Equation (III-C). We refer to this approach as block-diagonal CMA and leave a formal proof of this claim for future work. Importantly, this formulation allows the CMA update to be implemented in a sequential way, ensuring linear computational complexity with respect to the time horizon $T$.
Remark 7.
MPPI with block-diagonal CMA is the same as $\text{PI}^{2}$-CMA [13] without temporal averaging.
III-D Numerical experiments
We introduce four TO benchmark problems based on Hydrax [33]: Cartpole, DoubleCartPole, PushT and Humanoid. Control bounds are enforced with a penalty term in the cost function. In addition, for the PushT and Humanoid tasks, we rely on a zero-order spline interpolation to reduce the dimension of the problem. All results are averaged over six seeds. More details about the cost functions and additional visualizations are available online  [^1] https://github.com/ajordana/zoo-rob.
We compare the performance of Predictive Sampling, randomized smoothing, MPPI and MPPI-CMA. All algorithms use $2048$ samples per iteration. For MPPI and MPPI-CMA, we use a temperature of $\lambda=0.1$. For the CMA update, we use a full covariance as we did not observe that this was a computational bottleneck. Nevertheless, we provide results with block-diagonal CMA in Appendix E. Furthermore, we find that MPPI-CMA performs significantly better when using a separate step size for the mean update and the covariance update. For completeness, results using a shared step size are provided in Appendix E.
Figure 2 shows the cost evolution across iterations for each test problem. The performance of randomized smoothing varies across test problems and is highly sensitive to the step size choice. As expected, MPPI-CMA consistently outperforms MPPI. Interestingly, Predictive Sampling performs well given its simplicity.
(a) Cartpole
(b) DoubleCartPole
(c) PushT
(d) Humanoid
Figure 2: Cost according to the number of iterations for different test problems. The solid lines represent the median taken over six seeds.
IV Policy optimization
One of the main limitations of TO is that it is not always straightforward to use it to derive a controller. Indeed, performing MPC implies solving TO problems online, which can be extremely challenging due to limited compute time. Even though recent works were able to achieve impressive zero-order MPC demonstrations, these performances have not yet matched those obtained by RL. The main strength of RL is that it relies solely on policy evaluation at runtime. All the computation is moved offline where the goal is to search for a policy that will maximize performance across all possible initial conditions. This can be written:
$F(\theta)=\mathbb{E}\left[J(\theta,s)\right]$ (29)
where the expectation is taken over initial states $s$ and where $J(\theta,s)$ is the cost of the trajectory with initial state $s$ and with the controller $\pi_{\theta}$:
$J(\theta,s)$ $=\sum_{k=0}^{\infty}\gamma^{k}r(s_{k},a_{k})$ (30)
[…]
$\theta$ denote the parameters of the policy (e.g. the weights of a neural network). $r$ is the reward, $a$ is the control action. Note that we follow the RL formalism by trading the minimization of a cost function for the maximization of a reward. Furthermore, in contrast to TO, the optimization for each trajectory is done with an infinite horizon and a discount factor $\gamma$. We do not consider stochastic dynamics for simplicity as most robotic models are deterministic (e.g. locomotion or manipulation). However, the ideas presented in this section should extend to the stochastic dynamics case.
Remark 8.
The formulation in Equation (29) can easily be extended to encompass domain randomization by taking the expectation across multiple parameters of the environment.
Because the policy is deterministic, RL algorithms aiming to solve this formulation are called Deterministic Policy Gradient (DPG) algorithms [34]. While stochastic policies can be relevant in the context of games or partially observable settings, most robotic applications such as locomotion or manipulation can be solved with deterministic policies. In fact, in practice, even works that rely on RL algorithms derived with stochastic policies generally deploy a deterministic policy on the robot.
There exist two main sampling approaches to maximize (29). The first one is to directly rely on the zero-order techniques which implies sampling in the parameter space of a neural network, similarly to [35]. The second approach, used in RL, is to leverage the structure of the problem to sample in the action space. In this work, we are interested in the latter one. In the deterministic setting, the gradient can be estimated with the DPG theorem [34]:
$\partial_{\theta}J(\theta,s_{0})$ $=\sum_{k=0}^{\infty}\gamma^{k}\partial_{\theta}\pi(s_{k})\partial_{a}Q(s_{k},a_{k})$ (31)
where the $Q$-function is defined as
$Q(s,a)=\sum_{k=0}^{\infty}\gamma^{k}r(s_{k},a_{k})$ (32)
[…]
Consequently, we have
$\nabla_{\theta}F(\theta)$ $=\mathbb{E}\left[\sum_{k=0}^{\infty}\gamma^{k}\partial_{\theta}\pi_{\theta}(s_{k})\partial_{a}Q(s_{k},a_{k})\right]$ (33)
Note that we follow the formalism of [36] and express the policy gradient as an expectation over entire trajectories rather than making the $\gamma$-discounted state distribution explicit. Unfortunately, this formula cannot be directly implemented in an algorithm because of the presence of the $Q$-function. Indeed, the $Q$-function cannot be computed analytically. One solution is to estimate it with function approximation at each gradient descent step. Algorithms relying on this mechanism are called actor-critics. The actor refers to the policy while the critic refers to the $Q$-function. In practice, the $Q$-function is learned in a self supervised way via the Bellman equation. There exist many techniques to estimate the critic efficiently between gradient updates of the actor; we refer the reader to [11] for a comprehensive treatment of the subject.
Remark 9.
A key distinction between TO and policy optimization lies in the objective function: the policy optimization objective is averaged across various initial conditions. As a result, the objective cannot be computed analytically and must instead be estimated through sampling. Consequently, gradient estimates are inherently noisy and introduce stochasticity into the optimization process [37]. In light of our discussion on Langevin dynamics, this inherent stochasticity of RL algorithms could explain their empirical ability to avoid poor local minima.
IV-A Gaussian smoothing on the $Q$-function
DDPG [38] originally proposed to learn the $Q$-function with a neural network and directly differentiate it to obtain $\partial_{a}Q(s_{k},a_{k})$. This idea led to state-of-the-art RL algorithms such as TD3 [39]. In light of the success of randomized smoothing, a natural idea could be to use Gaussian smoothing to estimate the gradient of the Q-function and use the following estimate for the gradient of the cost:
$\mathbb{E}\left[\sum_{k=0}^{\infty}\gamma^{k}\partial_{\theta}\pi_{\theta}(s_{k})(Q(s_{k},a_{k}+\epsilon_{k})-Q(s_{k},a_{k}))\Sigma^{-1}\epsilon_{k}\right]$ (34)
where the expectation is taken with respect to both initial condition and the Gaussian random variables $\epsilon_{k}$. This provides us with a simple variation of actor-critics based on the DPG theorem. Algorithm (7) summarizes the procedure.
Input: $s_{0},\theta,Q$
1 while stopping criterion is not met do
2       Compute action $a=\pi_{\theta}(s)$
3       Sample $\epsilon$
4       Update the actor $\theta\leftarrow\theta+\alpha(Q(s,a+\epsilon)-Q(s,a))\partial_{\theta}\pi_{\theta}(s)\Sigma^{-1}\epsilon$
5       Update the critic, $Q$
6
Output: $\theta$
Algorithm 7 Randomized smoothing - Actor-Critic
We deliberately omit the rollout logic as well as the interplay between the actor update and the critic update. This is because our focus is solely on modifying the actor’s gradient update in DPG algorithms. We do not aim to provide a novel actor-critic logic. As we will see in the experimental section, this simple change can lead to competitive algorithms. An interesting variation of Algorithm (7) is to use the log-sum-exp transform and use exponential average weights in the actor update.
Remark 10.
One way to understand this algorithm is to interpret the partial derivative of the policy $\partial_{\theta}\pi_{\theta}(s)$ as a linear mapping from action space to the policy parameter space. Consequently, the algorithm can be understood as a way to map samples from action space into samples from the policy parameter space.
IV-B Numerical experiments
In this section, we investigate the benefits of estimating the gradient of the $Q$-function in DPG algorithms such as DDPG and TD3. Our implementation relies on CleanRL [40] and only modifies the actor’s update rule. Specifically, we investigate two smoothing approaches, the default randomized smoothing (RS) and the one relying on the log-sum-exp (LSE) transform. Consequently, we benchmark DDPG and TD3 against their smoothed counterparts: RS-DDPG, LSE-DDPG, RS-TD3, LSE-TD3. The performances are evaluated on seven Mujuco environments. For the sampling, we use 10 samples per update, with sampling noise following a Gaussian distribution with a standard deviation of $0.1$ (the covariance matrix is proportional to the identity matrix). Our implementation is available online  [^2] https://github.com/ajordana/zoo-rob.
Five runs are performed for each test problem. Figure 3 presents the normalized score across all runs. The episodic return for each environment is provided in Appendix F. The results indicate that both RS-DDPG and LSE-DDPG significantly outperform DDPG, which demonstrates the benefits of smoothing. For TD3, the improvements are not as clear. This could be explained by the fact that TD3 is already a very strong algorithm and that the margin of improvement on those benchmarks is limited. Additionally, our modifications were restricted to the actor’s update; achieving state-of-the-art performance would most likely require extensive hyperparameter tuning of the actor-critic mechanism. Nevertheless, the results show that deterministic policy gradient algorithms can benefit from smoothing.
Figure 3: Performance of each RL algorithm.
IV-C Connection to the Reinforce algorithm
One may ask how the update rule introduced in Equation (34) compares to the classical Reinforce algorithm. First, let’s recall how this algorithm is derived. In its most general form, Reinforce relies on the gradient of the following formulation.
$J(\theta,s_{0})=\mathbb{E}\left[\sum_{k=0}^{\infty}\gamma^{k}r_{k}(S_{k},A_{k})\right]$ (35)
[…]
The policy is no longer a deterministic function but a parametrized distribution over the space of possible actions. Note that the expectation is now taken over both the initial conditions and random actions. We now use capitalized letter for actions and states to emphasize their stochastic nature. The stochastic policy gradient theorem states that
$\frac{\partial}{\partial\theta}J(\theta,s_{0})$ $=\mathbb{E}\left[\sum_{k=0}^{\infty}\gamma^{k}Q(S_{k},A_{k})\frac{\partial}{\partial\theta}\log(p_{\theta}(A_{k}|S_{k}))\right]$
Note that the definition of the $Q$-function is no longer the same as in the deterministic case. It is now defined as the expectation of stochastic rollouts starting at state $s$ with action $a$. In practice, in most RL algorithms, the actions are sampled from a Gaussian probability distribution with mean $\pi_{\theta}(s)$ and covariance $\Sigma$ where $\pi_{\theta}$ is a neural network. Hence, we have:
$a\sim\mathcal{N}(\pi_{\theta}(s),\Sigma)$ (36)
Therefore,
$\frac{\partial}{\partial\theta}\log(p_{\theta}(a|s))=\frac{\partial\pi_{\theta}(s)}{\partial\theta}^{\top}\Sigma^{-1}(a-\pi_{\theta}(s))$ (37)
Consequently, we have:
$~{}\frac{\partial}{\partial\theta}J(\theta,s_{0})=$ $\mathbb{E}\left[\sum_{k=0}^{\infty}\gamma^{k}\left(Q(S_{k},A_{k})-V(S_{k})\right)\right.$
$\left.\frac{\partial\pi_{\theta}(S_{k})}{\partial\theta}^{\top}\Sigma^{-1}(A_{k}-\pi_{\theta}(S_{k}))\right]$ (38)
The baseline $V(S_{k})$ can be subtracted because the random variable $A_{k}-\pi_{\theta}(S_{k})$ has zero-mean conditioned on the state. Similarly to the randomized smoothing case this induces an invariance to translation by a constant of the reward function. Interestingly, we recover a formula very similar to Equation (34). The only difference is that the rollouts are now stochastic. This changes two things. The first one is that the $Q$-function and the partial derivatives are now evaluated at $S_{k}$ (instead of $s_{k}$ in Equation (34)). The second is that the definition of the $Q$-function has changed. In Equation (34), the $Q$-function denoted deterministic rollouts. In contrast, it is now an expectation of the outcome of stochastic rollouts.
The additional stochasticity in Equation (IV-C) (as compared to Equation (34)) can be understood as exploration noise. One formal justification could be to understand the stochastic policy in Equation (35) as a smoothing operating in action space. Similarly to Gaussian smoothing, the smoothing allows to derive the gradient of the cost with only $Q$-function evaluation. Consequently, the policy gradient theorem is analogous to the log-likelihood trick of Gaussian smoothing.
In the end, Reinforce can be seen as a way to solve the deterministic formulation introduced in Equation (29) without the simulator’s gradients by sampling in the action space. Interestingly, the smoothing of the Reinforce formulation introduces additional stochasticity as compared to Equation (34). It would be interesting to understand whether this extra stochasticity that can be interpreted as exploration noise is useful in practice. Intuitively, smoothing helps to simplify the problem but adding too much noise might render the surrogate too different from the initial problem. Future work could explore connections to other popular stochastic policy gradient algorithms such as TRPO or PPO.
V Population based algorithms
Theoretically, sampling-based algorithms following Langevin dynamics can sample from any probability distribution [23]. However, when applied to challenging multimodal probability distributions, the samples may fail to mix well and can exhibit strong autocorrelation. A natural way to address this issue is to explore algorithms that consider multiple starting points.
So far, all the algorithms that we have seen, either in TO or RL, have been local search techniques based on the simple greedy local search from Algorithm 2. It would be worth deriving algorithms combining local search concepts with global ones.
The most naive way to combine local and global search is to perform multiple local searches with diverse initial starting points. This is also known as Random Restart. Algorithm 8 summarizes the procedure.
Input: $x_{0}\in\mathbb{R}^{n}$
1 $x\leftarrow x_{0}$
2 while stopping criterion is not met do
3       Sample $\bar{x}$ in $\mathbb{R}^{n}$
4       $\tilde{x}\leftarrow\text{Greedy Local Search}(\bar{x})$
5       if $f(\tilde{x})<f(x)$  then
6             $x\leftarrow\tilde{x}$
7
8
Output: $x$
Algorithm 8 Random Restarts
A more sophisticated approach is to perform several local search in parallel. One way to do this is to maintain a population of $N$ samples, generate a new population by applying local search and then select the $N$ best samples. This procedure, called $(N+\lambda)$-ES, is one of the most popular ES algorithms. Algorithm 9 summarizes this procedure. Interestingly, there are two special cases: (1 + 1)-ES is equivalent to Greedy Local Search (Algorithm 2) and (1 + $\lambda$)-ES is equivalent to Predictive sampling [9] (Algorithm 4).
Input: Initial population of $N$ values: $x_{1},\dots,x_{N}$
1 while Stopping criterion is not met do
2       $D\leftarrow\{x_{1},\dots,x_{N}\}$
3       for $i=1,\dots\lambda$ do
4             $x\leftarrow$ Sample uniformly from $\{x_{1},\dots,x_{N}\}$
5             $d\leftarrow$ Sample a centered Gaussian
6             Add $x+d$ to $D$
7
8      $\{x_{1},\dots,x_{N}\}\leftarrow$ the $N$ elements in $D$ with smallest $f(x)$
9
Output: $x$
Algorithm 9 $(N+\lambda)$-ES
A potential limitation of this type of approach is that many greedy local searches may converge to the same local minima. Ideally, samples should coordinate with each other. An intuitive idea would be to avoid having particles too close to each other. One way to do so is to perform Stein Variational Gradient Descent (SVGD) [41]. This formulation naturally encompasses a repulsion term to maintain a certain distance between particles. Although SVGD requires gradients, it is natural to use the concepts covered in the previous Sections to derive a gradient free variation. For instance, [42] derives a gradient free SVGD based on CMA-ES.
VI Conclusion
In this work, we have demonstrated how random search provides a unifying perspective on zero-order algorithms commonly used in robotics. We also discussed theoretical concepts that help explain why sampling-based zero-order techniques can escape local minima. Leveraging this understanding, we proposed novel and competitive RL algorithms. These algorithms are only examples of the potential outcomes enabled by this unified viewpoint. In the future, we hope this tutorial will help the robotics community tackle open challenges such as constrained zero-order optimization and the search for global solutions.
Appendix A Connection between the greedy local search and the Metropolis-Hastings algorithm
[26] shows how the Metropolis adjusted Langevin algorithm can be interpreted as a line search with the Armijo condition. Similarly, in the trivial case where the step size $\alpha_{k}=0$, in Equation (14), we can recover a connection to Algorithm 2. The idea of the Metropolis-Hastings algorithm [23] is to accept a step from $x$ to $x^{\prime}$ with probability:
$\min\left(1,\frac{p(x^{\prime})}{p(x)}\frac{q(x|x^{\prime})}{q(x^{\prime}|x)}\right)$ (39)
If $\alpha_{k}=0$ in Equation (14), then,
$q(\cdot|x)\propto\mathcal{N}(x,\gamma^{2}I)$ (40)
Given Equation (12), the acceptance probability of the Metropolis-Hastings algorithm [23] is
$\min\left(1,\frac{p(x^{\prime})}{p(x)}\frac{q(x|x^{\prime})}{q(x^{\prime}|x)}\right)=\min\left(1,e^{-\left(f(x^{\prime})-f(x)\right)}\right)$
Interestingly, $e^{-\left(f(x^{\prime})-f(x)\right)}>1$ implies $f(x^{\prime})<f(x)$. Hence, one could easily transform the Greedy Local Search (Algorithm 2) in an MCMC method with the Metropolis-Hastings algorithm[23]. Indeed, in Algorithm 2, we could accept the step $x^{\prime}=x+d$ with probability one in case of decrease and allow for an increase with probability $e^{-\left(f(x^{\prime})-f(x)\right)}$.
Appendix B About the natural gradient
The idea of the natural gradient [27] is to find the steepest descent according to an arbitrary metric $d$. More precisely, the natural gradient looks for a small step $h$ satisfying:
$\min_{h}$ $f(x)+\nabla f(x)^{T}h$ (41)
s.t. $d(x+h,x)=\epsilon^{2}$
where $\epsilon$ is a small scalar. Typically, $d$ is such that (at least locally):
$d(x_{1},x_{2})=\frac{1}{2}(x_{2}-x_{1})^{T}A(x_{2}-x_{1})$ (42)
Let’s write the Lagrangian of the problem (41)
$\mathcal{L}(h,\lambda)=f(x)+\nabla f(x)^{T}h-\lambda(\frac{1}{2}h^{T}Ah-\epsilon^{2})$ (43)
We have:
$\nabla_{h}\mathcal{L}(h,\lambda)=0$ $\iff\nabla f(x)-\lambda Ah=0$
$\iff h=\frac{1}{\lambda}A^{-1}\nabla f(x)$ (44)
The sign of $\lambda$ can be recovered with the second order optimality condition [43].
$\nabla^{2}_{hh}\mathcal{L}(h,\lambda)\succeq 0\implies-\lambda A\succeq 0\implies\lambda\leq 0$ (45)
Hence, the natural gradient step is of the form:
$h\propto-A^{-1}\nabla f(x)$ (46)
In the context where $\theta$ characterizes a probability distribution, a typical choice of distance is:
$\operatorname{KL}(p_{\theta+h}||p_{\theta})=\int_{x}p_{\theta+h}(x)\log\left(\frac{p_{\theta+h}(x)}{p_{\theta}(x)}\right)dx$ (47)
Locally, we can show that this metric is quadratic. Let’s look at the partial derivatives of Equation (47) with respect to $h$.
$\frac{\partial\operatorname{KL}(p_{\theta+h}||p_{\theta})}{\partial h}=\int_{x}\partial_{h}p_{\theta+h}(x)\log\left(\frac{p_{\theta+h}(x)}{p_{\theta}(x)}\right)dx$
[…]
$\frac{\partial\operatorname{KL}(p_{\theta+h}||p_{\theta})}{\partial h}(h=0)=0$ (48)
Then, we can compute the Hessian with respect to $h$.
$\frac{\partial^{2}\operatorname{KL}(p_{\theta+h}||p_{\theta})}{\partial^{2}h}=\int_{x}\partial^{2}_{h}p_{\theta+h}(x)\log\left(\frac{p_{\theta+h}(x)}{p_{\theta}(x)}\right)dx$
$\;\;\;\;\;\;\;\;+\int_{x}\partial_{h}p_{\theta+h}(x)^{T}\partial_{h}p_{\theta+h}(x)\frac{1}{p_{\theta+h}(x)}dx$ (49)
And, we find that:
$\frac{\partial^{2}\operatorname{KL}(p_{\theta+h}||p_{\theta})}{\partial^{2}h}(h=0)$ (50)
$=\int_{x}\partial_{h}p_{\theta}(x)^{T}\partial_{h}p_{\theta}(x)\frac{1}{p_{\theta}(x)}dx$
$=\int_{x}\partial_{\theta}\log(p_{\theta}(x))^{T}\partial_{\theta}\log(p_{\theta}(x))p_{\theta}(x)dx$
$=\mathbb{E}_{X\sim p_{\theta}}\left[\partial_{\theta}\log(p_{\theta}(X))^{T}\partial_{\theta}\log(p_{\theta}(X))\right]=F(\theta)$
where $F(\theta)$ is the Fisher matrix. In the end,
$\operatorname{KL}(p_{\theta+h}||p_{\theta})=\frac{1}{2}h^{T}F(\theta)h+o(||h||^{2})$ (51)
Let’s consider the specific case where $\theta=m$ and:
$p_{\theta}(x)=\frac{1}{\kappa}\exp\left(-\frac{1}{2}(x-m)^{T}\Sigma^{-1}(x-m)\right)$
We have
$\partial_{\theta}\log(p_{\theta}(x))=(x-\theta)^{T}\Sigma^{-1}$ (52)
Therefore, the Fisher matrix reads:
$F(\theta)$ $=\mathbb{E}\left[\Sigma^{-1}(x-\theta)(x-\theta)^{T}\Sigma^{-1}\right]$ (53)
$=\Sigma^{-1}\mathbb{E}\left[(x-\theta)(x-\theta)^{T}\ \right]\Sigma^{-1}$
$=\Sigma^{-1}$
Hence, in that case, the Fisher matrix is equal to the inverse of the covariance matrix. The case where $\theta=(m,\Sigma)$ can be found in [31].
Appendix C About the connection between the log-sum-exp and risk-sensitive control
Interestingly, the surrogate defined in Equation (18) reminds of the risk-seeking control formulation [29]. At first, this can seem counter-intuitive as in practice, we prefer designing risk-averse controllers rather than risk-seeking ones. However, from an optimization perspective, the risk-seeking formulation has the advantage of giving more importance to lower values. Figure 4 illustrates the landscape for each formulation. The risk-neutral formulation corresponds to the default randomized smoothing. One way to interpret these landscapes is to see the smoothing noise as a disturbance from the environment. On one hand, the risk-seeking formulation expects the noise to act in its favor by shifting the state towards the global solution. Therefore, it will lower the function value in the surroundings of the narrow global minima. On the other hand, the risk-averse formulation expects the noise to be adversarial. Therefore, it will increase the function value (as compared to the neutral formulation) around the narrow global minima.
Figure 4: Risk-seeking and Risk-averse transformation
Appendix D Connection to the Cross Entropy Method (CEM)
The Cross Entropy Method (CEM) is a popular algorithm that can be used for optimization. Algorithm 10 summarizes the main step of CEM. Interestingly, this algorithm connects to the natural gradient [30]. Indeed, in Equation (III-C), one can assign $w_{k}=\frac{1}{K_{e}}$ to the best $K_{e}$ samples and 00 to the others (by best, we mean samples with the lowest value $f(x_{k})$). Choosing a step-size $\alpha=1$ yields an algorithm close to the well-known Cross-Entropy Method (CEM). The only difference with the original CEM is that natural gradient update computes the covariance update with the old $x$ (i.e. the one used to sample) instead of the new one [30].
Input: $x_{0}\in\mathcal{X},\Sigma_{0},K_{e}$
1 $(x,\Sigma)\leftarrow(x_{0},\Sigma_{0})$
2 while stopping criterion is not met do
       Sample $x_{1:K}\sim\mathcal{N}(x,\Sigma)$
        // sample
       Sort $x_{1:K}$ such that $f(x_{1})\leq f(x_{2})\leq\dots\leq f(x_{K})$
        // Rank
       $x\leftarrow\sum\limits_{k=1}^{K_{e}}\frac{1}{K_{e}}x_{k}$
        // update
       $\Sigma\leftarrow\sum\limits_{k=1}^{K_{e}}\frac{1}{K_{e}}(x_{k}-x)(x_{k}-x)^{\top}$
        // update
3
Output: $x$
Algorithm 10 Cross-Entropy Method (CEM)
Appendix E Additional TO results
Figure 2 provides additional results for the TO benchmark. Specifically, we show how block-diagonal CMA compares to the default CMA (relying on the full matrix). In addition, we illustrate the advantage of having different step-sizes for the mean and the covariance.
(a) Cartpole
(b) DoubleCartPole
(c) PushT
(d) Humanoid
Figure 5: Cost according to the number of iterations for different test problems. The solid line represents the median taken over 6 seeds.
Appendix F Additional RL results
Figure 6 presents an alternative metric summarizing the RL results across tasks (instead of runs). Figure 7 shows the episodic return for each test environment.
Figure 6: Performance of each RL algorithm. The metric is evaluated per task.
Figure 7: Episodic return for each environment